OpenStudy (anonymous):

7x+5y=6 3x+y=6 solve for xy

OpenStudy (anonymous):

7x+5y=6 3x+y=6 -> y = 6-3x Sub 6-3x for y 7x+5(6-3x)=6 7x+30-15x=6 -8x+30=6 -8x=-24 x=3 3(3)+y=6 9+y=6 y=-3